OpenStudy (anonymous):

r = <9, -7, -1>, v = <2, 2, -2>, w = <-5, -2, 6> v ⋅ w

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

\[ \mathbf v = <\color{red}{2}, \color{blue}2, \color{green}{-2}>, \mathbf w = <\color{red}{-5}, \color{blue}{-2}, \color{green}6> \]\[ \mathbf v\cdot \mathbf w = \color{red}{2\times -5}+ \color{blue}{2\times -2}+\color{green}{-2\times 6} \]

OpenStudy (anonymous):

-10-4-12=-28

OpenStudy (anonymous):

*-26

OpenStudy (anonymous):

You can do it on your own.